Give the slope of any line perpendicular to the given line.
Y= -2/7x+2
slope: m= -2/7
y-intercept: b= 2

Answer by ewatrrr(24785)

Hi,
.Per your reply
Y= -4/-3x-8 Slope: -4/-3x Y-intercept= -8
always works the same way: the slope of any line perpendicular to this line would be the negative reciprocal of the slope given for this line.
.
In this case m = (-4/-3)or m = (4/3)
.
the negative reciprocal of that would be (-3/4)
.
Hi,
*Note: The slope of a line perpendicular to a given line: is equal to the negative reciprocal, of the slope of the line given.
.
In this case:
y = -2/7x + 2
.
Any line perpendicular to it would have a slope of 7/2